# Knowledge Accumulation in Continually Learned Representations and the Issue of Feature Forgetting

### v3

@@ Abstract
Continual learning research has shown that neural networks suffer from catastrophic forgetting “at the output level”, but it is debated whether this is also the case at the level of learned representations. Multiple recent studies ascribe representations a certain level of innate robustness against forgetting - that they only forget minimally and no critical information. We revisit and expand upon the experiments that revealed this difference in forgetting and illustrate the coexistence of two phenomena that affect the quality of continually learned representations: knowledge accumulation and feature forgetting. Carefully taking both aspects into account, we show that, even though it is true that feature forgetting can be small in absolute terms, newly learned information tends to be forgotten just as catastrophically at the level of the representation as it is at the output level. Next we show that this feature forgetting is problematic as it substantially slows down knowledge accumulation. Finally, we study how feature forgetting and knowledge accumulation are affected by different types of continual learning methods.
@@ 1 Introduction
Machine learning models are typically trained on static datasets, and once they are deployed, they are usually not updated anymore. However, sometimes models make mistakes. Sometimes they do not work in a domain that was not trained. Sometimes they do not recognize certain classes or corner cases. To overcome such malfunctions, the default choice in industry is to gather new data and to retrain a model from the beginning with both the new and old data (Huyen, 2022). Retraining a full model is costly and time-consuming, especially in deep learning. The goal of continual learning is to enable models to train continually, to learn from new data when they become available. This has proven to be a hard challenge (De Lange et al., 2022; van de Ven et al., 2022), as deep learning models that are continually trained exhibit catastrophic forgetting (McCloskey & Cohen, 1989; Ratcliff, 1990). Without precautionary measures, new information is learned at the expense of forgetting earlier acquired knowledge.
@@
In continual learning, there are at least two good reasons to care about representations. First, a strong representation makes it easier to learn new information (Kornblith et al., 2019). When a model already has a good representation, it may require less changes to adapt to new data, which lowers the level of inference and thus the risk of forgetting (Cha et al., 2021). Second, progressively accumulating knowledge from individual tasks into one representation is a goal on its own. Effective continual learning should be able to use new information to its benefit and build a stronger representation over time (Chen & Liu, 2018), which can finally be used to solve a variety of tasks (Bengio et al., 2013).
It is with these motivations that recent works have been studying how representations are learned in continual learning, and how they forget. Multiple studies observe an apparent robustness to forgetting for representations (Davari et al., 2022; Zhang et al., 2022; Hu et al., 2022; Wang et al., 2023). Davari et al. (2022) write: “[…] in many commonly studied cases of catastrophic forgetting, the representations under naive fine-tuning approaches, undergo minimal forgetting, without losing critical task information” (p. 16713), and Zhang et al. (2022) write: “there seems to be no catastrophic forgetting in terms of representations” (p. 4) and “common techniques for mitigating catastrophic forgetting […] have little effect on improving [representations]” (p. 4). When measuring the representation quality for a task that was not included in the training sequence, Zhang et al. (2022) further conclude that “representation learning and catastrophic forgetting are largely separate issues” (p. 9). These studies suggest that when forgetting of learned features happens, it does not hinder learning good representations. If this is true, forgetting would only be a problem if one cares about the performance on the trained tasks, but not if you care about learning a good general representation.
@@
To determine the impact of feature forgetting on the quality of the continually learned representation for downstream tasks, we compare the representation of a continually trained model against a representation that is ensembled from a model copy obtained after task one, after task two, and so on. This ensemble baseline and the continually trained model learn in the exact same way, but differ because the former does not forget. We find that the ensemble baseline yields a substantially better general representation than the continually trained model, which indicates that feature forgetting negatively affects knowledge accumulation. This means that preventing feature forgetting is not only important for the performance on tasks that a model was trained on, but also to learn strong representations in general.
Most experiments in this paper study the continual learning and forgetting of representations in supervised learning, though in Appendix C we show evidence that suggests our main conclusions also hold for self-supervised learning. We conclude the paper by evaluating examples of important families of continual learning methods and report how they influence the learning and forgetting of representations.
In summary, our contributions include [^1] Code will be made public upon acceptance.:
  • We illustrate the characteristics of feature forgetting and knowledge accumulation evaluated at the representation level (Section 3)
@@
  • We show that forgetting of the representation hinders knowledge accumulation (Section 4.3).
@@ 2 Preliminaries
We follow the common definition of a continual learning setting by assuming a sequence of $m$ disjoint tasks $\mathcal{T}=\{T_{1},T_{2},...,T_{m}\}$. Each task consists of training data with inputs $X_{i}$ and targets $Y_{i}$, as well as respective test data $\hat{X}_{i}$, $\hat{Y}_{i}$. While training on task $T_{j}$, only the training data $\{X_{j},Y_{j}\}$ of that task are used. Exceptions are replay memories, which can store small subsets of data from past tasks; we only use such replay memories when studying experience replay in Section 5. Finally, we use $T_{d}$ to refer to a downstream task, i.e. one that is not part of the training sequence $\mathcal{T}$. To study the representation learned by a model on the sequence $\mathcal{T}$, we distinguish between the backbone parameters $\theta_{B}$ and the task specific heads $\theta_{H}=\{\theta_{h_{1}},...,\theta_{h_{m}}\}$. For task $T_{j}$, the output of a model is calculated as $g(f(x;\theta_{B});\theta_{H})$ and its representation as $f(x;\theta_{B})$. In the remainder of this paper the heads $\theta_{h_{j}}$ are linear layers and we use, with slight abuse of notation, $f(x;{\theta_{h_{j}},\theta_{B}})$ to refer to the combination of backbone and head. Alternatively, we use $f_{i}$ to denote the entire model after training task $T_{i}$ and $f_{0}$ for a network with random weights.
@@
Our main focus is on classification tasks. We use $r^{f}_{i,j}$ as a general symbol to refer to a metric that quantifies the performance or quality of model $f_{i}$ for task $T_{j}$. We drop super- and subscript when they are clear from context. To differentiate between the quality of the backbone and the entire model, we define two such performance metrics: output accuracy $\text{OUT}_{i,j}$ and linear probe accuracy $\text{LP}_{i,j}$ (or LP accuracy). $\text{OUT}_{i,j}$ refers to the standard test accuracy of model $f(x;\theta_{h_{j}},\theta_{B})$ immediately after training task $T_{i}$. To measure the quality of a representation, we would ideally know the optimal head $\theta^{*}_{h_{j}}$ for a particular backbone $\theta_{B}$ and task $T_{j}$. To approximate this optimal head at test time, we train a new head with parameters $\tilde{\theta}_{h_{j}}$ using the training data of task $T_{j}$ while keeping the backbone parameters $\theta_{B}$ frozen, as is common in representation learning (Bengio et al., 2013). Using this new head, $\text{LP}_{i,j}$ is equal to the test accuracy of $f(x;\tilde{\theta}_{h_{j}},\theta_{B})$ on task $T_{j}$ immediately after training task $T_{i}$. Alternative ways to evaluate representations reach similar results for the experiments reported in this paper, see Appendix G.
@@
In the main paper, the reported results are on Split MiniImageNet, a 20 task (five classes each) split of MiniImageNet (Vinyals et al., 2016). We use 19 tasks as the training sequence $\mathcal{T}$, while the remaining task is never seen during training and used as downstream task $T_{d}$. To reduce the influence of the difficulty of a task, we use five task splits and report mean and standard error (SE) on all results. The splits are randomly generated but consistent across experiments. Further details of our experimental protocol are reported in Appendix A. In Appendix F, we replicate our experiments using Split CIFAR-100 (Krizhevsky et al., 2009).
@@ 3 Knowledge accumulation and feature forgetting
Catastrophic forgetting is observed when measuring accuracy at the output level of a continually finetuned neural network. However, recent studies exemplify how evaluating performance of the same model at representation level instead leads to a different result. In the following, we highlight two evaluation perspectives that are used in recent literature to illustrate this difference between the output and representation level. With these evaluation perspectives, two phenomena can be identified that affect the quality of continually learned representations: feature forgetting and knowledge accumulation.
@@
Figure 1: Illustration of feature forgetting and knowledge accumulation, for a model continually finetuned on Split MiniImageNet. On the left, the output accuracy and LP accuracy of $T_{6}$ are shown. The difference between LP accuracy directly after the task was trained and the LP accuracy after training $T_{i}$ is what we call feature forgetting. On the right, the same models are evaluated with the same metrics but on a downstream task ($T_{d}$). We refer to the improvement in LP accuracy over that of a randomly initialized model ($f_{0}$) as knowledge accumulation. (Mean $\pm$ SE, 5 runs)
The evaluation perspectives differ in what data they use. On the one hand, a task that was included in the training sequence (e.g. $T_{6}$) can be used, on the other hand, a downstream task ($T_{d}$) that is excluded from the training sequence can be used. Each perspective can show one of the two phenomena. Using e.g. $T_{6}$, it can be observed that linear probing, i.e. retraining the head, can recover much of the performance that was lost, which could be interpreted as representations forgetting less than at the output (Davari et al., 2022). The performance degradation, or forgetting, that persists beyond the recovered performance of the linear probe is what we call feature forgetting, see Figure 1 (left). The second evaluation perspective inspects the performance on a downstream task using the same models $f_{i}$ as the previous setting. The representation for $T_{d}$ progressively improves with each learned task, without significant declines. We call these successive improvements knowledge accumulation, see Figure 1 (right). Zhang et al. (2022); Cha et al. (2022); Yoon et al. (2023) had observed this upward trend and concluded that feature forgetting does not pose an issue for learning representations continually. In the next sections, we continue to use both concepts to revisit the dynamics of continual learning at the representation level.
@@ 4.1 Relative forgetting
McCloskey & Cohen (1989) and Ratcliff (1990) are often credited for discovering the phenomenon of catastrophic forgetting. They describe it respectively as: “[t]raining on a new set of items may drastically disrupt performance on previously learned items” (p. 110), and “well-learned information is forgotten rapidly as new information is learned” (p. 285). In recent literature, following Lopez-Paz & Ranzato (2017), forgetting is commonly quantified as the difference between the performance immediately after task $T_{j}$ was trained and the performance after $n$ new tasks: $r_{j,j}-r_{j+n,j}$. Both output accuracy and linear probing accuracy have been used in this way (e.g. Davari et al., 2022). Forgetting, defined as such, is thus equal to the absolute drop in performance. In contrast, the amount of forgetting can also be considered relative to how much information was learned during training the task. We propose to define the relative forgetting of task $T_{j}$ after $n$ new tasks as:
@@
Or, in words: relative forgetting measures how much of the performance that was gained from training on a task is forgotten afterwards. Figure 3 shows why this matters. When inspecting the absolute forgetting at the output and the representation (middle panel), one could conclude that the forgetting at the output is worse than at the representation level. However, the representation quality also improved less during training of task $T_{6}$ (left panel), partly because of previously accumulated knowledge. Relative to the gained performance, forgetting is roughly equal at the representation and at the output (right panel); both lose nearly all gained performance. Considering the gained performance is thus crucial when comparing forgetting in settings where the performance before training a task differs across the compared settings, because it changes the denominator of Equation 1. This happens when considering representations and LP accuracy, but in most cases not for output accuracy. Their difference is comparable to the difference between the answers to the following two questions before seeing any task data: Which test samples belong to the unknown category x? and Given that x looks like this, which other test samples are of category x?. While the first answer will be random, the second one depends on how good the description, e.g. the representation, of x is. A model that has accumulated more knowledge and has a better representation can have improved performance on the second question, but not on the first.
@@
Figure 3 shows the relative forgetting of the representation and of the output for tasks $1$ until $9$ (the one of $T_{6}$ is thus a copy of the right panel in Figure 3). For early tasks (especially $T_{1}$), forgetting as measured by linear probing stabilizes after one or two new tasks, while for the output $\text{rF}_{n,j}$ continues to increase. For the later tasks, the representation forgets at least as much as at the output (see Appendix E for task $T_{10}$ and more). This observation noticeably leads to a different conclusion than the one drawn by the papers in the introduction: except for the first tasks, almost all performance that was gained while training a task, is lost when training on new ones, both at the output and the representation level.
@@
Figure 2: Performance, absolute forgetting and relative forgetting of $T_{6}$ for a model continually finetuned on Split MiniImageNet. Comparing output accuracy and representation quality (LP accuracy) can have different interpretations when considered in absolute or relative terms. Absolute forgetting (middle) suggests that output forgetting is worse than at the representation level. But, it does not account for performance already accumulated before learning $T_{6}$ (left). Expressing forgetting relative to newly gained performance (right) reveals that forgetting trends are similar at the output and representation level. (Mean $\pm$ SE, 5 runs)
Figure 2: Performance, absolute forgetting and relative forgetting of $T_{6}$ for a model continually finetuned on Split MiniImageNet. Comparing output accuracy and representation quality (LP accuracy) can have different interpretations when considered in absolute or relative terms. Absolute forgetting (middle) suggests that output forgetting is worse than at the representation level. But, it does not account for performance already accumulated before learning $T_{6}$ (left). Expressing forgetting relative to newly gained performance (right) reveals that forgetting trends are similar at the output and representation level. (Mean $\pm$ SE, 5 runs)
@@
Figure 3: Relative forgetting at the output level and at the level of representations, as per Eq. 1, for a model continually fine-tuned on the Split MiniImageNet sequence. Relative to the knowledge gained during training on the task, forgetting in the representation and at the output are similar. (Mean $\pm$ SE, 5 runs)

### v4

@@ Abstract
Continual learning research has shown that neural networks suffer from catastrophic forgetting “at the output level”, but it is debated whether this is also the case at the level of learned representations. Multiple recent studies ascribe representations a certain level of innate robustness against forgetting – that they only forget minimally in comparison with forgetting at the output level. We revisit and expand upon the experiments that revealed this difference in forgetting and illustrate the coexistence of two phenomena that affect the quality of continually learned representations: knowledge accumulation and feature forgetting. Taking both aspects into account, we show that, even though forgetting in the representation (i.e. feature forgetting) can be small in absolute terms, when measuring relative to how much was learned during a task, forgetting in the representation tends to be just as catastrophic as forgetting at the output level. Next we show that this feature forgetting is problematic as it substantially slows down the incremental learning of good general representations (i.e. knowledge accumulation). Finally, we study how feature forgetting and knowledge accumulation are affected by different types of continual learning methods.
@@ 1 Introduction
Machine learning models are typically trained on static datasets, and once they are deployed, they are usually not updated anymore. However, sometimes models make mistakes. Sometimes they do not work in a domain that was not trained on. Sometimes they do not recognize certain classes or corner cases. To overcome such malfunctions, the default choice in industry is to gather new data and to retrain a model from the beginning with both the new and old data (Huyen, 2022). Retraining a full model is costly and time-consuming, especially in deep learning. The goal of continual learning is to enable models to train continually, to learn from new data when they become available. This has proven to be a hard challenge (De Lange et al., 2022; van de Ven et al., 2022), as deep learning models that are continually trained exhibit catastrophic forgetting (McCloskey & Cohen, 1989; Ratcliff, 1990). Without precautionary measures, new information is learned at the expense of forgetting earlier acquired knowledge.
@@
In continual learning, there are at least two good reasons to care about representations. First, a strong representation makes it easier to learn new information (Kornblith et al., 2019). When a model already has a good representation, it may require less changes to adapt to new data, which lowers the risk of forgetting (Cha et al., 2021). Second, continually learning a good representation by progressively accumulating knowledge from individual tasks is a goal on its own. Effective continual learning should be able to use new information to its benefit and build a stronger representation over time (Chen & Liu, 2018), which can finally be used to solve a variety of tasks (Bengio et al., 2013).
It is with these motivations that recent works have been studying how representations are learned in continual learning, and how they forget. Multiple studies observe an apparent robustness to forgetting for representations (Davari et al., 2022; Zhang et al., 2022; Hu et al., 2022; Wang et al., 2023; Anthes et al., 2023). For example, Davari et al. (2022) write: “[…] in many commonly studied cases of catastrophic forgetting, the representations under naive fine-tuning approaches, undergo minimal forgetting, without losing critical task information” (p. 16713), and Zhang et al. (2022) comment: “there seems to be no catastrophic forgetting in terms of representations” (p. 4) and “common techniques for mitigating catastrophic forgetting […] have little effect on improving [representations]” (p. 4). These studies further suggest that in the case that some forgetting of learned features does happen, this does not hinder the learning of good general representations. For example, when measuring the representation quality for a task that was not included in the training sequence, Zhang et al. (2022) conclude that “representation learning and catastrophic forgetting are largely separate issues” (p. 9). If this is true, forgetting would only be a problem if one cares about the performance on the trained tasks, but not if one cares about learning a good general representation.
@@
To determine the impact of feature forgetting on the quality of the continually learned representation for downstream tasks, we compare the representation of a continually trained model against a representation that is ensembled from copies of the model obtained after finishing training on each task. This ensemble baseline and the continually trained model learn in the exact same way, but differ because the former does not forget. We find that the ensemble baseline yields a substantially better general representation than the continually trained model, which indicates that feature forgetting slows down knowledge accumulation (i.e. the improvement of the learned representation over time). This means that preventing feature forgetting is not only important for the performance on tasks that a model was trained on, but also to learn strong representations in general.
Most experiments in this paper study the continual learning and forgetting of representations in supervised learning, though in Appendix C we show evidence that suggests that our main conclusions also hold for self-supervised learning. We conclude the paper by evaluating examples of important families of continual learning methods and reporting how they influence the learning and forgetting of representations.
In summary, our contributions include [^1] Code available at: https://github.com/TimmHess/KAaFF:
  • We illustrate the characteristics of feature forgetting and knowledge accumulation evaluated at the representation level (Section 3).
@@
  • We show that forgetting in the representation hinders knowledge accumulation (Section 4.3).
@@ 2 Preliminaries
We follow the common definition of a continual learning setting by assuming a sequence of $m$ disjoint tasks $\mathcal{T}=\{T_{1},T_{2},...,T_{m}\}$. Each task $T_{j}$ consists of training data with inputs $X_{j}$ and targets $Y_{j}$, as well as respective test data $\hat{X}_{j}$, $\hat{Y}_{j}$. While training on task $T_{j}$, only the training data $\{X_{j},Y_{j}\}$ of that task are used. Exceptions are replay memories, which can store small subsets of data from past tasks; we only use such replay memories when studying experience replay in Section 5. Finally, we use $T_{d}$ to refer to a downstream task, i.e. one that is not part of the training sequence $\mathcal{T}$. To study the representation learned by a model on the sequence $\mathcal{T}$, we distinguish between the backbone parameters $\theta_{B}$ and the task-specific heads $\theta_{H}=\{\theta_{h_{1}},...,\theta_{h_{m}}\}$. For task $T_{j}$, the output of a model is calculated as $g(f(x;\theta_{B});\theta_{h_{j}})$ and its representation as $f(x;\theta_{B})$. In this paper, each head $\theta_{h_{j}}$ specifies a linear layer and we use, with slight abuse of notation, $f(x;{\theta_{h_{j}},\theta_{B}})$ to refer to the combination of backbone and head. Alternatively, we use $f_{i}$ to denote the entire model directly after training on task $T_{i}$, and we use $f_{0}$ for a model with random weights.
@@
Our main focus is on classification tasks. We use $r^{f}_{i,j}$ as a general symbol to refer to a metric that quantifies the performance or quality of model $f_{i}$ for task $T_{j}$. We drop super- and subscript when they are clear from context. To differentiate between the quality of the backbone and the entire model, we define two such performance metrics: output accuracy $\text{OUT}_{i,j}$ and linear probe accuracy $\text{LP}_{i,j}$ (or LP accuracy). $\text{OUT}_{i,j}$ refers to the standard test accuracy of model $f(x;\theta_{h_{j}},\theta_{B})$ immediately after training task $T_{i}$. When $i<j$, output accuracy is equal to random performance, i.e. the performance of randomly initialized head. To measure the quality of a representation, one option would be to use the optimal head $\theta^{*}_{h_{j}}$ for a particular backbone $\theta_{B}$ and task $T_{j}$. To approximate this optimal head, we train a new head with parameters $\tilde{\theta}_{h_{j}}$ using the training data of task $T_{j}$ while keeping the backbone parameters $\theta_{B}$ frozen, as is common in representation learning (Bengio et al., 2013). Using this new head, $\text{LP}_{i,j}$ is equal to the test accuracy of $f(x;\tilde{\theta}_{h_{j}},\theta_{B})$ on task $T_{j}$, whereby $\theta_{B}$ are the backbone parameters immediately after training on task $T_{i}$. Alternative ways to evaluate representations, such as using $k$NN, reach similar results for the experiments reported in this paper, see Appendix H.
@@
In the main paper, the reported results are on Split MiniImageNet, a 20 task (five classes each) split of MiniImageNet (Vinyals et al., 2016). We use 19 tasks as the training sequence $\mathcal{T}$, while the remaining task is never seen during training and used as downstream task $T_{d}$. To reduce the influence of the difficulty of a task, we use five task splits and report mean and standard error (SE) on all results. The splits are randomly generated but consistent across experiments. When evaluating the performance on a task, the model only needs to choose between the classes of that task. We do not include all classes seen so far (i.e., as would be done in a class-incremental learning evaluation), because this would make the evaluation ‘task’ more difficult after each trained task, thereby confounding the analysis. Further details of our experimental protocol are provided in Appendix A. In Appendix G, we report experiments using Split CIFAR-100 (Krizhevsky et al., 2009).
@@ 3 Knowledge accumulation and feature forgetting
Figure 1: Illustration of feature forgetting and knowledge accumulation, for a model continually finetuned on Split MiniImageNet. On the left, the output accuracy and LP accuracy on $T_{6}$ are shown. The difference between LP accuracy directly after the task was trained and the LP accuracy after training on later tasks is what we call feature forgetting. On the right, the same continually trained model is evaluated with the same metrics but on a downstream task ($T_{d}$). We refer to the improvement in LP accuracy over that of a randomly initialized model ($f_{0}$) as knowledge accumulation. (Mean $\pm$ SE, 5 runs)
Catastrophic forgetting is observed when measuring accuracy at the output level of a continually finetuned neural network. However, recent studies exemplify how evaluating performance of the same model at representation level instead leads to a different result. In the following, we highlight two evaluation perspectives that have been used in recent literature to illustrate this difference between the output and representation level. With these evaluation perspectives, two phenomena can be identified that affect the quality of continually learned representations: feature forgetting and knowledge accumulation. The evaluation perspectives differ in what data they evaluate on. On the one hand, a task that was included in the training sequence (e.g. $T_{6}$) can be used, on the other hand, a downstream task ($T_{d}$) that is excluded from the training sequence can be used. Each perspective can show one of the two phenomena.
Evaluating on a task included in the training sequence (e.g. $T_{6}$), it can be observed that linear probing, i.e. retraining the head, can recover much of the performance that is lost in the output accuracy (Figure 1, left), which could be interpreted as representations forgetting less than at the output (Davari et al., 2022). The performance degradation, or forgetting, that persists beyond the recovered performance of the linear probe is what we call feature forgetting. When features are forgotten, the representations of past classes become less separable. This contributes to the total amount of forgetting in continual learning, to which e.g. representation drift and misaligned heads contribute as well (Caccia et al., 2021), see also Section 7.
The second evaluation perspective inspects the performance of a continually trained model on a downstream task. Using the same setting as above, we find that the representation for $T_{d}$ progressively improves with each learned task, without significant declines (Figure 1, right). We call this improvement of a representation for unseen tasks knowledge accumulation. A model that has accumulated more knowledge by this definition, is a more general model and is better suited to solve new tasks, although it may have forgotten features specific to a past task at the same time. This definition is similar to the one used by Jin et al. (2021), who equate knowledge accumulation with the capability of generalizing to unseen tasks. Lesort et al. (2023) and Caccia et al. (2020) use the term knowledge accumulation slightly differently, to loosely refer to performance on both old and new tasks. Knowledge accumulation is related to forward transfer, which is one of the main desiderata in task-incremental learning (van de Ven et al., 2022; De Lange et al., 2022). Yet task-incremental learning results themselves are not well suited to study the representations directly, as representation drift (Caccia et al., 2021) can result in misaligned heads, which could suggest that there is no knowledge accumulation, even when there is. Finally, in continual pretraining (Hu et al., 2022; Lee et al., 2023; Cossu et al., 2022), the performance on downstream tasks is what matters, yet the influence of forgetting pretrained features has not been studied in these works. In what follows, we use both concepts, feature forgetting and knowledge accumulation, to study the dynamics of continual learning at the representation level.
@@ 4 Forgetting in continually learned representations
In this section, we analyze feature forgetting by taking knowledge accumulation into account. Section 4.1 shows that newly gained performance tends to be lost similarly at the output and the representation. Features learned during the first tasks in the training sequence are an exception to this, which we investigate further in Section 4.2. Finally, Section 4.3 points out that feature forgetting hinders effective knowledge accumulation for downstream tasks.
@@ 4.1 Relative forgetting
McCloskey & Cohen (1989) and Ratcliff (1990) are often credited for discovering the phenomenon of catastrophic forgetting. They describe it respectively as: “[t]raining on a new set of items may drastically disrupt performance on previously learned items” (p. 110), and “well-learned information is forgotten rapidly as new information is learned” (p. 285). In recent literature, following Lopez-Paz & Ranzato (2017), forgetting is commonly quantified as the difference between the performance immediately after task $T_{j}$ was trained on and the performance after training on $n$ new tasks: $r_{j,j}-r_{j+n,j}$. Both output accuracy and linear probing accuracy have been used in this way (e.g. Davari et al., 2022). Forgetting, defined as such, is thus equal to the absolute drop in performance. In contrast, the amount of forgetting can also be considered relative to how much information was learned during training on the task. We propose to define the relative forgetting of task $T_{j}$ after $n$ new tasks as:
@@
Or, in words: relative forgetting measures how much of the performance that was gained during training on a task is forgotten afterwards. Figure 3 illustrates how absolute forgetting and relative forgetting can lead to different conclusions. When inspecting the absolute forgetting at the output and the representation (middle panel), one could conclude that the forgetting at the output is worse than at the representation level. However, the representation quality also improved less during training of task $T_{6}$ (left panel), partly because of previously accumulated knowledge. Relative to the gained performance, forgetting is roughly equal at the representation and at the output (right panel); both lose nearly all gained performance. Considering the gained performance is thus crucial when comparing forgetting in settings where the performance before training a task differs across the compared settings, because it changes the denominator of Equation 1. This happens when considering representations and LP accuracy, but in most cases not for output accuracy. Their difference is comparable to the difference between the answers to the following two questions before seeing any task data: Which test samples belong to the unknown category x? and Given that x looks like this, which other test samples are of category x?. While the first answer will be random, the second one depends on how good the description, e.g. the representation, of x is. A model that has accumulated more knowledge and has a better representation can have improved performance on the second question, but not on the first.
@@
Figure 3 shows the relative forgetting of the representation and of the output for tasks $1$ until $9$ (the one of $T_{6}$ is thus a copy of the right panel in Figure 3). For early tasks (especially $T_{1}$), forgetting as measured by linear probing stabilizes after one or two new tasks, while for the output $\text{rF}_{n,j}$ continues to increase. For the later tasks, the representation forgets at least as much as at the output (see Appendix F for task $T_{10}$ and more). This observation noticeably leads to a different conclusion than the one drawn by the papers in the introduction: except for the first tasks, almost all performance that was gained while training a task, is lost when training on new ones, both at the output and the representation level.
@@
Figure 2: Performance, absolute forgetting and relative forgetting of $T_{6}$ for a model continually finetuned on Split MiniImageNet. Comparing output accuracy and representation quality (LP accuracy) can have different interpretations when considered in absolute or relative terms. Absolute forgetting (middle) suggests that output forgetting is worse than at the representation level. Such absolute forgettting does not account for performance already accumulated before learning $T_{6}$ (left). When expressing forgetting relative to newly gained performance (right), forgetting trends are similar at the output and representation level. (Mean $\pm$ SE, 5 runs)
@@
Figure 3: Relative forgetting at the output level and at the level of representations, as per Eq. 1, for a model continually fine-tuned on the Split MiniImageNet sequence. Relative to the knowledge gained during training on the task, forgetting in the representation and at the output are similar, except for the first few tasks. (Mean $\pm$ SE, 5 runs)